OpenStudy (anonymous):

The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool?

OpenStudy (campbell_st):

start with A = length x width so in your question \[7x^3 -42x^2 = 7x^2 \times length\] by changing the subject it becomes \[\frac{7x^2}{7x^2} - \frac{42x^2}{7x^2} = length\] simplify the fractions for the length of the pool

OpenStudy (anonymous):

So what did you get?

OpenStudy (campbell_st):

I haven't done it... I've just given you the process to do it

OpenStudy (anonymous):

Oh can you do it?

OpenStudy (anonymous):

Please???

OpenStudy (campbell_st):

well I'd look at the problem like this \[\frac{7x^2 \times x}{7x^2} - \frac{6 \times 7x^2}{7x^2} = length\] just simplify for the length, hope you can see the common factor

OpenStudy (anonymous):

so it would be....?

OpenStudy (anonymous):

??????????

OpenStudy (campbell_st):

well have an attempt at simplifying the fractions...its reasonably obvious